Reports 1-1 of 1 Clear search Modify search
AdV-ALS (Green sources)
gosselin - 19:08 Thursday 23 March 2023 (59397) Print this report
ALS WEB fiber arrangement and green conversion efficiency

I noticed that the correction of the PSTAB at WE were oscillating quite a lot since Tuesday. This is symptomatic of a lack of power reaching the ALS elec box. As a consequence, sometimes the 2W amplifier was not seeded with enough power and its internal securtiy was switching it off (plot 1) 
We know that the fibers are quite sensitive, for what concern the power reaching the elec box. We suspected that the problem was coming from some displacement of the fiber during the acivities at WE on Tuesday.
I went there and by rearranging the fiber I managed to get back the corrections to the level close to what they were (lower than 2V, see plot 2)

I took advantage of being there to check the temperature set point of the oven of the PPLN crystal used to convert the IR in green. I slightly changed the setpoint and checked the conversion efficiency: GREEN / IR^2. 
The setpoint was looking good but at some point, while doing the tests the conversion efficiency suddently increased (see plot 3). This looks too fast to be due to a temperature change in the oven (less than 1s). It is not explained and quite strange. It happened two more times in the following hour (plot 4). 
We will monitor the behavior in the coming days and investigate the phenomena. 

Images attached to this report
Comments to this report:
gosselin - 10:03 Friday 24 March 2023 (59406) Print this report

The sudden changes in the power seen by WEB_PD_GREEN_MONI_CALI are coincident with the opening/closing of the flip mirror inside the SHGB (see plot 1).
The actual power reaching this monitoring photodiode is quite low (about 150 uW), the changes are most likely due to some diffused light coming from the beam dump when the mirror is flipped. 
The problem is not present at NEB (plot 2)

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×